Homep Page Num Lock Page 7 8 9 PgUp 6 4. 5 2) A coil of 160 turns and area 0.20 m2 is placed with its axis parallel to a magnetic field 2 of initial magnitude 0.40 T. The magnetic field changes uniformly from 0.40 T in the +x direction to 0.40 T in the -x direction in 2.0 s. If the resistance of the coil is 16 ?, at what rte is power generated in the coil? A) B)
Answer #1

Here ,

area , A = 0.20 m^2

for the current in the coil

I = change in magnetic flux/(time * resistance)

I = 160 * 0.20 * (0.40 - (-0.40))/(2 * 16)

I = 0.8 A

the induced current is 0.8 A

rate of power generated = I^2 * R

rate of power generated = 0.8^2 * 16

rate of power generated = 10.24 W